Michael has a ladder set up to clean the gutters of his house. The top of the ladder reaches the base of the gutter, 30 feet from the ground. If the ladder is 32 feet tall, how far from the base of the house is the base of the ladder?


1
Expert's answer
2021-12-06T16:09:22-0500

"d=\sqrt{32^2-30^2}=\sqrt{124}\approx 11.14\;ft."